Find the probability that 3 or 4 cars have been sold in a day



"\begin{array}{c:c}
 x &Pr(X=x) \\
 \hdashline
0 & 0.25
\\ 
 \hdashline
1 & 0.2\\
 \hdashline
2 & 0.15\\
 \hdashline
3 & 0.1\\
 \hdashline
 4 & 0.15
\\
 \hdashline
 5 & 0.15
\end{array}"
1
Expert's answer
2021-06-29T09:50:39-0400
"P(X=3\ or\ X=4)=P(X=3)+P(X=4)"

"=0.1+0.15=0.25"

The probability that 3 or 4 cars have been sold in a day is 0.25.